OpenStudy (hippoharry):

Given the data points (1, 4), (4, 13), and (6, 19), which of the following equations describes the best-fit line? A. y = 1 + 3x B. y = 3 + 1x C. y = -1 + 3x D. y = 2 + 3x

OpenStudy (peachpi):

plug the points into the lines to see which one all 3 points fits

OpenStudy (emyoung0):

take two of the points and find the gradient, it will be the same even if you take the third point. e.g 13-4/4-1 = 3 eliminate all options without a 3x. Now use the point gradient formula: y-y1 = m(x-x1) so pick any of the three, im picking no 1.. y-4 = 3(x-1) expand and simplify so y= 1 + 3x

OpenStudy (hippoharry):

oh i get it i understand. so all i had to do was plug in each and see which on has all the points

OpenStudy (peachpi):

yes